OpenStudy (anonymous):

I need to find the surface area of a cylinder in terms of pi with a radius of 4 cm and a height of 5cm.

OpenStudy (zzr0ck3r):

cylinder has surface area \( \large S=2\pi*r*h+2\pi*r^2\)

OpenStudy (anonymous):

okay so I get to \[2\pi20+2\pi16\] but I am not sure what I multiply together to get the final answer

OpenStudy (zzr0ck3r):

that is the final answer

OpenStudy (zzr0ck3r):

\(2\pi20+2\pi16=2\pi(20+16)=2\pi*36=72\pi\)

OpenStudy (anonymous):

So 72pi is the final answer? Thank you so much!

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

np